# ToDA: Target-oriented Diffusion Attacker against Recommendation System

### v1

@@
Abstract.
Recommendation systems (RS) have become indispensable tools for web services to address information overload, thus enhancing user experiences and bolstering platforms’ revenues. However, with their increasing ubiquity, security concerns have also emerged. As the public accessibility of RS, they are susceptible to specific malicious attacks where adversaries can manipulate user profiles, leading to biased recommendations, aka. shilling attacks. Recent research often integrates additional modules using generative models to craft these deceptive user profiles, ensuring them are imperceptible while causing the intended harm. Albeit their efficacy, these models face challenges of unstable training and the exploration-exploitation dilemma, which can lead to suboptimal results.
In this paper, we pioneer to investigate the potential of diffusion models (DMs), for shilling attacks. DMs have showcased remarkable capabilities in areas like image synthesis, recommendation systems, and adversarial attacks, providing finer control over the generation process. Specifically, we propose a novel Target-oriented Diffusion Attack model (ToDA). It incorporates a pre-trained autoencoder that transforms user profiles into a high dimensional space, paired with a Latent Diffusion Attacker (LDA)—the core component of ToDA. LDA introduces noise into the profiles within this latent space, adeptly steering the approximation towards targeted items through cross-attention mechanisms. The global horizon, implemented by a bipartite graph, is involved in LDA and derived from the encoded user profile feature. This makes LDA possible to extend the generation outwards the on-processing user feature itself, and bridges the gap between diffused user features and target item features. Extensive experiments compared to several SOTA baselines demonstrate ToDA’s effectiveness. Specific studies exploit the elaborative design of ToDA and underscore the potency of advanced generative models in such contexts.
@@ Abstract.
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXX
1. Introduction
@@
Recommendation system primarily endeavors to capture users’ preferences through their historical interactions, thereby predicting potential item candidates that would likely attract users (Rendle et al., 2012; Koren, 2008; He et al., 2020). As an effective countermeasure against information overload in a web system, it remarkably enhances the user experience while concurrently boosting the revenue of merchants in many web services (e.g., e-commerce (Ma et al., 2022) and content-sharing platform (Wei et al., 2019)).
Figure 1. An illustrative example of shilling attacks on recommendation system. A manually crafted malicious user $u_{4}$ and its associate interactions would make the recommendation results biased to certain intended item $i_{5}$.
@@
In light of the ubiquitous nature and public accessibility of recommendation systems (i.e., recommendation models are learned based on user data which is usually accessible), the necessity for prioritizing security has become increasingly apparent (Fan et al., [n. d.]; Deldjoo et al., 2021). Although many works leverage adversarial learning with disturbed samples to improve the robustness (Du et al., 2018; Yuan et al., 2019; Tang et al., 2019), recommendation system is still vulnerable to malicious attacks (Zeng et al., 2023; Lin et al., [n. d.], 2022; Nguyen Thanh et al., 2023; Zhang et al., [n. d.]; Cohen et al., 2021). In other words, attackers are able to glean interaction histories of users and subsequently construct fabricated user profiles (i.e., a sequence of user-item interactions) as inputs to the recommendation system, thereby promoting or demoting the target items. This action is recognized as shilling attack. Figure 1 depicts the example of a shilling attack. The attacker crafts a fake user $u_{4}$ who has interacted with three items, $i_{3}$, $i_{4}$, and $i_{5}$ where $i_{5}$ is the target item. With the training of historical and injected interactions, recommendation system updates its parameters, resulting in a prediction offset. Specifically, the target items gain a higher probability of being candidates in the recommendation list. Obviously, shilling attacks undermine the integrity of recommendation systems, leading to an unfair exposure of items, and consequently eroding users’ trust.
@@
Several efforts are devoted to the design of shilling attacks as the security concerns and insights provided to defense-side research. Initial works primarily depend on heuristic rules to manually craft user profiles, like Random Attack (Lam and Riedl, 2004) and Bandwagon Attack (Gunes et al., 2014), and approximate an optimization problem (Li et al., 2016; Fang et al., 2018; Huang et al., [n. d.]). However, generative models gradually take the dominant role of shilling attacks to the present. These methods can be categorized into autoencoder (AE)-based (Zeng et al., 2023; Zhang et al., [n. d.]), generative adversarial network (GAN)-based (Wu et al., 2021; Wang et al., 2022; Lin et al., [n. d.], 2022; Nguyen Thanh et al., 2023), and reinforcement learning (RL)-based (Chiang et al., 2023; Zhang et al., 2022; Chen et al., [n. d.]; Fan et al., 2021; Zhang et al., 2023a; Song et al., [n. d.]) strategies, which are in an autoagressive manner. To adapt these methods for shilling attack tasks, an additional module is often incorporated to optimize the attack objectives (Zhang et al., [n. d.]; Lin et al., 2022; Wu et al., 2021). Despite their efficacy, these models face challenges such as unstable training and the exploration-exploitation dilemma, which can lead to suboptimal results. Clearly, the role of the generative model in shilling attack strategies remains an area ripe for further exploration to fully harness its potential.
@@
Diffusion models (DMs) have emerged as a cutting-edge technique for generating data across various domains, like computer vision (Croitoru et al., 2023) and natural language processing (Reid et al., 2022; Gong et al., 2022). At their core, these models simulate the process of diffusion, denoising the data step by step to reconstruct the data. Some works in recommendation system (RS) harness DMs to predict user preference under the noisy scenario (Wang et al., 2023b) and forecast the users’ tastes evolving over time (Li et al., 2023). In adversarial attack field, DMs helps to purify samples’ perturbations (Nie et al., 2022), craft malicious visual samples (Chen et al., 2023) and improve model robustness (Wang et al., 2023a; Xiao et al., 2022). DMs provide finer control over the generation process, allowing for enhanced interpretability and precision. Their inherent denoising capability also ensures the generation of high-quality data. Given the success of DMs in these areas, there is sparked interest in their potential to shilling attacks, an intriguing area where DMs remain largely unexplored.
@@
Introducing DMs into shilling attack tasks presents unique challenges. First, the inherent nature of DMs is benign. They are designed to understand and replicate patterns without any malicious intent. Even though some attempts in RS and adversarial attack showcase the potential, tailoring DMs specifically for shilling attacks requires significant modifications to ensure the generated profiles are not only imperceptible but also effective in manipulating the victim RS. Second, DMs conventionally have a narrow focus, usually concentrating on a single sample during the generation process. In contrast, shilling attacks require a global view that outwards the on-processing pattern. This is essential to identify beneficial items that can be used as auxiliaries to ensure the fake user profiles are both imperceptible and malicious.
@@
To address the above challenges, we propose a novel Target-oriented Diffusion Attack model, termed ToDA. ToDA hinges on a latent diffusion model by first encoding user profiles into a high-dimensional space to diminish the computation cost and facilitate model flexibility. Therefore, we are able to subtly add noise to latent features at each forward step, and employ an approximator to reconstruct every state, referencing the prior step during the reversion process. This procedure allows us to sample natural profiles that conform to victim RS, inherently ensuring the desired imperceptibility. To steer the latent diffusion model from being malicious, we harness the Target-oriented Approximator within the reversion by using cross-attention to condition the target item’s features. Without an extra module or attack objective, we endow the diffusion model with attack ability. Through a bipartite graph (i.e., user-item graph), we are able to generate target item features in a global view. This makes it possible to extend the generation of a broader horizon outwards the on-processing user feature itself, and bridges the gap between diffused user features and target item features. Wherein, a SOTA GNN encoder in collaborative filtering (i.e., LightGNN (He et al., 2020)) is adopted to catch the high-order correlations of target items. Our ToDA is simple yet flexible and powerful for shilling attacks. We compare ToDA with several SOTA shilling attack models, like LegUP (Lin et al., 2022) and GSPAttack (Nguyen Thanh et al., 2023), and exploit significant improvements. Overall, our contributions are threefold:
@@
  • We devise a novel target-oriented diffusion attacker, ToDA, underpinned by the latent diffusion model paired with a target-oriented approximator. We innovatively confers attack ability to DMs, filling a blank in the confluence of DMs and shilling attacks.
@@
2. Preliminary
We introduce the basic procedure of shilling attacks and determine the capability of attackers. The concept of recommendation system is also formally involved.
Shilling Attacks: Given the gleaned interactions $\mathbf{Y}\in\{0,1\}^{n\times m}$, the attacker crafts a set of malicious user profiles, where $n$ and $m$ denote the number of users and items, respectively. We denote these shilling users as $\mathcal{U}^{a}=\{u^{a}_{1},u^{a}_{2},\dots,u^{a}_{k}\}$, where $k=|\mathcal{U}^{a}|$. The corresponding interaction matrix for these users is represented as $\mathbf{Y}^{a}\in\{0,1\}^{k\times m}$. Formally, we represent this procedure as $\mathbf{Y}^{a}=A(\mathbf{Y},\mathcal{T},k)$, where $A(\cdot)$ is the shilling attack algorithm that takes the gleaned interactions as input and outputs the $k$ fake user profiles. And $\mathcal{T}$ denotes the indices of target items. Note that, $\mathbf{Y}$ is sparser than the integrated interactions due to the capability of attackers and the constraint of recommendation systems. Finally, the attacker injects these malicious user profiles into the recommendation system. This effectively extends the original user-item matrix to $\mathbf{Y}^{\prime}\in\{0,1\}^{(n+k)\times m}$, which includes both the original and the crafted user-item interactions. The capability of attackers is determined from three perspectives:
@@ 2. Preliminary
  • Incomplete data: Shilling attackers often operate with limited information, reflecting only a subset of the total user-item interactions. Consequently, the malicious user profiles, $\mathcal{U}^{a}$, and corresponding interaction matrix, $\mathbf{Y}^{a}$, are built based on this incomplete data set.
@@
  • Black-box setting: Attackers typically do not have access to the inner workings of the recommendation system, including its learning algorithm and parameter settings. This is often referred to as a black-box setting. The shilling attack algorithm, $A(\cdot)$, therefore, operates blindly, producing $\mathbf{Y}^{a}$ without detailed insights about the system it’s intended to manipulate.
@@
  • No extra knowledge: The third capability implies that the attacker doesn’t possess additional knowledge outside of the gleaned interactions, $\mathbf{Y}^{\prime}$. This includes knowledge about item features, user demographics, or other auxiliary information that could enhance the attack’s precision.
Recommendation System: Here we denote the user set as $\mathcal{U}=\{u_{1},u_{2},\dots,u_{n}\}$ and the item set =$\mathcal{I}=\{i_{1},i_{2},\dots,i_{m}\}$. The interactions between users and items are encapsulated in a matrix $\tilde{\mathbf{Y}}\in\{0,1\}^{n\times m}$, where the value in the row $u$ and column $i$ signifies an interaction between user $u$ and item $i$. The chief goal of recommendation system is to compute the preference score $\hat{\mathbf{Y}}_{u,i}$ indicating the potential interest of user $u$ in item $i$. More formally, the receommendation system $R$ harnesses collaborative filtering signals to derive this as $\hat{\mathbf{Y}}_{u,i}=R(u,i,\tilde{\mathbf{Y}})$.
3. Methodology
By exhibiting the potential of DMs and addressing the mentioned challenges, we introduce the crux of ToDA, the Latent Diffusion Attacker, adding noise to latent features and reversing them through Target-oriented Approximator, and the training and inference of ToDA.
3.1. Latent Diffusion Attacker

### v3

@@
Abstract
Recommendation system (RS) has become an indispensable tool to address information overload, simultaneously enhancing user experiences and bolstering platforms’ revenues. However, due to the public accessibility, it is susceptible to specific malicious attacks where attackers can manipulate user profiles, leading to biased recommendations, aka. shilling attacks. Recent research uses generative models and integrates additional modules to craft deceptive user profiles, ensuring they are imperceptible while causing the intended harm. Despite the effectiveness, these models face challenges of learning dilemmas and inflexibility, which can lead to suboptimal performance.
In this paper, we propose a novel Target-oriented Diffusion Attack model (ToDA), pioneering the investigation of the potential of diffusion models (DMs). DMs have showcased remarkable capabilities in areas like image synthesis, recommendation systems, and adversarial attacks, providing finer control over the generation process. To assimilate DMs within shilling attacks, we address their inherent benign nature and the narrowness of the local view. ToDA incorporates a pre-trained autoencoder that transforms user profiles into a high-dimensional space, paired with a Latent Diffusion Attacker (LDA). LDA introduces noise into the profiles within the latent space, adeptly steering the approximation towards targeted items through cross-attention mechanisms. The global view, implemented by a bipartite graph, enables LDA to extend the generation beyond the on-processing user feature itself and bridges the gap between diffused user features and target item features. Extensive experiments compared to several SOTA baselines demonstrate ToDA’s efficiency and efficacy, highlighting its potential in both DMs and shilling attacks.
Index Terms:
@@ Abstract
I Introduction
@@
Recommendation system (RS) primarily endeavors to capture users’ preferences through their historical interactions, thereby predicting potential item candidates that would likely attract users [1, 2, 3]. As an effective countermeasure against information overload, it remarkably enhances the user experience while concurrently boosting the revenue of merchants in many web services (e.g., e-commerce [4] and content-sharing platform [5]).
Figure 1: Illustration of a shilling attack example and the comparison between conventional diffusion models and ToDA, which derives the targeted profile by leveraging global view graph and target information.
@@
In light of the ubiquitous nature and public accessibility of recommendation systems (i.e., recommendation models are trained based on user data which is usually accessible), the necessity for prioritizing security has become increasingly apparent [6, 7]. Although many works leverage adversarial learning with disturbed samples to improve the robustness [8, 9, 10], recommendation system is still vulnerable to malicious attacks [11, 12, 13, 14, 15, 16, 17]. In other words, attackers are able to glean interaction histories of users and subsequently construct fabricated user profiles (i.e., a sequence of user-item interactions) as inputs to the recommendation system, thereby promoting or demoting the target items. This action is recognized as shilling attack. Figure 1 (a) depicts the example of a shilling attack. The attacker crafts a fake user $u_{4}$ who has interacted with three items, $i_{3}$, $i_{4}$, and $i_{5}$ where $i_{5}$ is the target item. With the training of historical and injected interactions, recommendation system updates its parameters, resulting in a prediction shift. Specifically, the target items gain a higher probability of being candidates in the recommendation list. Obviously, shilling attacks undermine the integrity of recommendation systems, leading to unfair exposure of items, and consequently eroding users’ trust.
@@
Several efforts are devoted to the design of shilling attacks as the security concerns and insights provided to defense-side research. The field of shilling attack research has seen a clear evolution. Initially, 1) researchers primarily adopt heuristic rules to manually craft user profiles [18, 19] or approximate an optimization problem [20, 21, 22]. Afterward, 2) generative models gradually take the dominant role of shilling attacks to the present: these methods can be categorized into autoencoder (AE)-based [11, 16], generative adversarial network (GAN)-based [23, 24, 12, 13, 14, 15], and reinforcement learning (RL)-based [25, 26, 27, 28, 29, 30] strategies, which are in an autoagressive manner. To adapt these methods for shilling attack tasks, an additional module is often incorporated to optimize the attack objectives [16, 14, 23]. Notably, recent shilling attack methods have been entrenched in GANs or autoregressive RL methods for several years since 2020 [12, 13]. Despite their efficacy, these models face challenges, such as unstable training and the exploration-exploitation dilemma, which lead to suboptimal results. To break down the standstill of recent shilling attacks and mitigating the inflexibility and learning dilemmas, a new generative paradigm for future shilling attacks is necessary.
@@
Diffusion models (DMs) have emerged as a cutting-edge technique for generating data across various domains, like computer vision [31] and natural language processing [32, 33]. At their core, these models simulate the process of diffusion, denoising the data step by step to reconstruct the data. Some works in recommendation system (RS) harness DMs to predict user preference under the noisy scenario [34] and forecast the users’ tastes evolving over time [35]. In adversarial attack field, DMs help to purify samples’ perturbations [36], craft malicious visual samples [37] and improve model robustness [38, 39]. DMs provide finer control over the generation process, allowing for enhanced interpretability and precision. Their inherent denoising capability also ensures the generation of high-quality data. Given the success of DMs in these areas, there is growing interest in their potential for shilling attacks, an intriguing area where DMs remain largely unexplored.
Utilizing DMs is natural yet challenging, presenting unique challenges. As illustrated in Figure 1 (b), we summarize the following two main challenges.
@@
  • The inherent nature of DMs is benign. They are designed to understand and replicate patterns without any malicious intent. Even though some attempts in RS and adversarial attack showcase the potential, tailoring DMs specifically for shilling attacks requires significant modifications to ensure the generated profiles are both imperceptible and effective in manipulating the victim RS.
  • DMs typically have a narrow focus, often concentrating on a single sample during generation (i.e., local view). However, shilling attacks necessitate a broader perspective that outwards the on-processing pattern to recognize auxiliary items (i.e., global view). For example, attackers must comprehend the interrelationships among various items and their mutual influences to identify a range of items relevant to the target, instead of merely assigning target items to the fabricated profiles.
To address the above challenges, we propose a novel Target-oriented Diffusion Attack model, termed ToDA. As shown in Figure 1 (c), ToDA derives targeted profiles step by step during the reverse procedure, where the global view graph and target information are incorporated.
@@
Specifically, ToDA hinges on a latent diffusion model by first encoding user profiles into a high-dimensional space to diminish the computational cost and facilitate model flexibility. Therefore, we are able to subtly add noise to latent features at each forward step, and employ an approximator to reconstruct every state, referencing the prior step during the reversion process. This procedure allows us to sample natural profiles that conform to victim RS, inherently ensuring the desired imperceptibility. To steer the latent diffusion model from being malicious, we harness the Target-oriented Approximator within the reversion by using cross-attention to condition the target item’s features under the global view. To this end, we adopt a bipartite graph (i.e., user-item graph), thus making it possible to extend the generation of a broader horizon outwards the on-processing user feature itself, and generating more diverse and relevant target item features. Wherein, a SOTA GNN encoder in collaborative filtering (i.e., LightGNN [3]) is adopted to catch the high-order correlations of target items. Without an extra module or attack objective, we endow the diffusion model with attack ability. Our ToDA is simple yet flexible and powerful for shilling attacks. We compare ToDA with several classical methods and SOTA generative models in the context of shilling attacks, like LegUP [14] and GSPAttack [15], and exploit significant improvements. Overall, our contributions are threefold:
@@
  • We devise a novel target-oriented diffusion attacker, ToDA, underpinned by the latent diffusion model paired with a target-oriented approximator. We innovatively confer attack ability to DMs, filling a blank in the confluence of DMs and shilling attacks.
@@
II Preliminary
In this section, we elucidate the goals of shilling attacks and determine the attacker capabilities in manipulating these systems, including incomplete data, black-box setting, and no extra knowledge. These settings ensure that our proposed attack are adapted to real-world scenarios. Furthermore, we provide a formal overview of recommendation system.
II-A Shilling Attacks
II-A1 Attack goal
In the context of recommendation systems, shilling attacks represent a deliberate effort to manipulate the recommendation process by injecting fraudulent user profiles into the system. These attacks aim to distort the recommendation outcomes in favor of certain items or to undermine the integrity of the recommendation algorithm. Formally, let $\mathbf{Y}\in\{0,1\}^{n\times m}$ denote the matrix of observed user-item interactions, where $n$ represents the number of users and $m$ denotes the number of items. The attacker crafts a set of malicious user profiles denoted as $\mathcal{U}^{a}=\{u^{a}_{1},u^{a}_{2},\dots,u^{a}_{k}\}$, with $k=|\mathcal{U}^{a}|$, and generates the corresponding interaction matrix $\mathbf{Y}^{a}\in\{0,1\}^{k\times m}$ for these fabricated users. This process is formalized as follows:
$\mathbf{Y}^{a}=A(\mathbf{Y},\mathcal{T},k),$ (1)
where $A(\cdot)$ represents the shilling attack algorithm, which takes the existing user-item interactions $\mathbf{Y}$, the set of target items $\mathcal{T}$, and the desired number of fake profiles $k$ as inputs, and produces the manipulated interaction matrix $\mathbf{Y}^{a}$. Subsequently, the attacker injects these malicious user profiles into the recommendation system, augmenting the original user-item matrix to $\mathbf{Y}^{\prime}\in\{0,1\}^{(n+k)\times m}$, thus influencing the recommendation outcomes. Following previous works [12, 13, 14, 15], our method ToDA focus on promotion attacks, which aim to maximize the availability of the target items in the recommendation list, in the current study to demonstrate the effectiveness and efficiency.
II-A2 Attacker capability
The efficacy of shilling attacks is also contingent upon various factors that delineate the attacker’s capabilities. These factors are assessed through a practical lens as outlined below:
@@ II-A2 Attacker capability
  • Incomplete data: Shilling attackers are assumed to operate with limited information, reflecting only a subset of the total user-item interactions. Prior works [12, 13, 14] often leverage the complete interactions of RS, countering the real scenario with system’s access constrains. Therefore, we define that the malicious user profiles, $\mathcal{U}^{a}$, and corresponding interaction matrix, $\mathbf{Y}^{a}$, are built upon this incomplete data set.
@@
  • Black-box setting: Attackers do not have access to the inner workings of the recommendation system, including its learning algorithm and parameter settings. The shilling attack algorithm, $A(\cdot)$, therefore, operates blindly, producing $\mathbf{Y}^{a}$ without detailed insights about the system. There is a general solution that introduce surrogate model trained with gleaned data to represent the victim model [14, 15]. However, such method is extensively time-consuming since it requires continuously updating the surrogate model. In contrast, ToDA can optimize the attack objective without surrogate model-aided and achieves outperformed efficacy than SOTA.
@@
  • No extra knowledge: Attacker doesn’t possess additional knowledge outside of the gleaned interactions, $\mathbf{Y}^{\prime}$. This includes knowledge about item features, user demographics, or other auxiliary information that could enhance the attack’s precision. Due to the incomplete data, several methods turn to introduce cross-domain data or knowledge graph to enhance the attack capability [28, 11], then boosting the attack performance, or exploring on domain-specific recommendations (like review-based [25] or sequential recommendation [40]). However, we aim to conduct the practical attack without extra knowledge against general recommendation system, then fundamentally facilitating the development of shilling attacks.
II-B Recommendation System
Let $\mathcal{U}=\{u_{1},u_{2},\dots,u_{n}\}$ represent the set of users and =$\mathcal{I}=\{i_{1},i_{2},\dots,i_{m}\}$ denote the set of items. The implicit interactions between users and items are encoded in a binary matrix $\tilde{\mathbf{Y}}\in\{0,1\}^{n\times m}$, where $\tilde{\mathbf{Y}}_{u,i}$ equals $1$ if user $u$ has interacted with item $i$, and 00 otherwise.
The primary objective of recommendation system is to predict the preference score $\hat{\mathbf{Y}}_{u,i}$, estimating the likelihood of user $u$ engaging with item $i$. This preference score is computed by a recommendation model $R$, which leverages collaborative filtering techniques to analyze the historical interactions captured in $\tilde{\mathbf{Y}}$. Mathematically, the recommendation model is represented as:
$\hat{\mathbf{Y}}_{u,i}=R(u,i,\tilde{\mathbf{Y}}),$ (2)
where $R(\cdot)$ utilizes the observed user-item interactions to infer the preference of user $u$ towards item $i$. Through this process, the recommendation system aims to enhance user satisfaction and engagement by delivering relevant and personalized recommendations.
III Methodology
We introduce the overall framework of our proposed method (in Section III-A), and specifically elucidate the crux of ToDA (as shown in Figure 2), the Latent Diffusion Attacker (in Section III-B), which adds noise to latent features and reverse them through Target-oriented Approximator (in Section III-C), and the optimization of ToDA (in Section III-D). Moreover, the complexity analysis showcases the efficiency of proposed method (in Section III-E).
Figure 2: The overall framework of ToDA. It is featured by the latent diffusion attacker with target-oriented approximator $\boldsymbol{\epsilon}_{\theta}$. The approximator $\boldsymbol{\epsilon}_{\theta}$ denoises the previous state of user profile feature $\mathbf{e}_{u;s}$, bringing in the global information to the local view diffusion model. The features of target items are incorporated through a GCN Encoder $\mathcal{E}_{\mathcal{G}}$, before using the cross-attention to control the denoising orientation.
III-A Approach Overview
Our goal is to introduce new paradigm of utilizing diffusion models to generate the user profile, achieving effective shilling attacks. Such paradigm avoids the previous learning dilemmas and inflexibility, opening a new door for recent still researches of shilling attacks. To this end, we point out several inevitable challenges, and establish a set of design principles that guide the development of our model:
  • Attack-Effectiveness (P1): A pivotal criterion in our framework is the effectiveness of the shilling attack simulation. This involves the generation of fabricated user profiles designed to manipulate the recommendation systems. To endow the malicious intent to conventional DMs, we aim to construct beneficial target item features, serve as the condition to guide the generation of DMs.
  • Generation-Efficiency (P2): Given the dynamic nature of recommendation systems and the computational intensity associated with conventional DMs, we recognize the need for a design that prioritizes efficiency in the generation process. Our approach incorporates sophisticated mechanisms to streamline the generation of synthetic profiles, thereby reducing the computational overhead and facilitating the scalability of both shilling attack simulations and DM applications.
  • Imperceptibility (P3): The imperceptibility of synthetic user profiles is crucial for the efficacy of shilling attacks. However, in the context of our research, this is considered a secondary objective. We acknowledge that while the ability to create imperceptible profiles is important, it is not the primary focus of our study. Nonetheless, the high-fidelity simulation capabilities inherent in DMs and GAN-based methods naturally contribute to the generation of profiles that are difficult to distinguish from genuine user data.
In alignment with the aforementioned principles, we introduce novel designs that not only enhance their utility in the context of shilling attack tasks but also address the limitations of conventional DMs:
  • Global Target-oriented Approximator (D1): To fulfill the Attack-Effectiveness (P1), we augment DMs with the capacity to simulate shilling attacks by integrating conditional generative frameworks. Our approach transcends the traditional localized interactions of DMs by employing a global interaction graph. This graph enables the generation of a comprehensive and informative target item representation that guides the generative process with distinct and robust signals, ensuring a higher fidelity in simulating targeted shilling attacks.
  • Lightweight Architecture (D2): Acknowledging the computational complexity of DMs, we propose a streamlined architecture (P2) for the dynamics of recommendation system. Our design eliminates redundant and cumbersome components, opting instead for a more efficient latent process that significantly reduces computational demands. Furthermore, we have replaced the traditional U-Net architecture with our proposed Target-oriented Approximator (ToA), as detailed in (D1). This novel architecture retains the core conceptual framework of DMs while simplifying complex and repetitive modules, leading to a more lightweight and agile model.
To integrate design principles with diffusion models, we introduce the Latent Diffusion Attacker (LDA), which is innated with a high-fidelity simulation capability (P3). Moreover, the LDA transcends the mere application of pre-existing diffusion models by incorporating a lightweight latent diffusion process, being complemented by a sophisticated mechanism for guiding the construction of user profiles, tailored to target items.
III-B Latent Diffusion Attacker